OpenStudy (anonymous):

How do I simplify this? 4^(log4(x+3))

OpenStudy (jdoe0001):

\(\bf \large {{\color{red}{ a}}^y={\color{blue}{ b}}\implies log_{\color{red}{ a}}{\color{blue}{ b}}=y \\ \quad \\ \quad \\ {\color{red}{ 4}}^{log_{\color{red}{ 4}}(x+3)}= \square \implies log_{\color{red}{ 4}}{\color{blue}{ \square }}=log_{\color{red}{ 4}}(x+3) }\) what do you think?

myininaya (myininaya):

\[a^{\log_a(f(x))}=f(x) , f(x)>0, a \in (0,1)\cup (1,\infty)\]

OpenStudy (jdoe0001):

or you can always keep in mind that \(\Large \bf \textit{log cancellation rule of }\quad {\color{red}{ a}}^{log_{\color{red}{ a}}x}=x\)

myininaya (myininaya):

since y=a^x and y=log_a(x) are inverses

OpenStudy (anonymous):

I still dont understand, can you provide an example please?

myininaya (myininaya):

example: \[a^{\log_a( \text{ candy} } ) =\text{ candy }\]

OpenStudy (anonymous):

ohh, I see. Thank you.

myininaya (myininaya):

so what do you think your answer is?

OpenStudy (anonymous):

X+3?

myininaya (myininaya):

yep

OpenStudy (anonymous):

Does this also work for log10^(10^3x+1)

OpenStudy (anonymous):

and the answer would be 3x+1?